Solve the system of equations by graphing. Then classify the system as consistent or inconsistent; and the equations as dependent or independent.
x+y=5
x-y=-1
What is the solution of the system of equations.
Please help me with this.

Answer by oscargut(2103)

x+y=5
x-y=-1
if you add equations
then
2x=4
x=2
then
2+y=5
y=3
solution is x=2,y=3 (unique solution so is independent)